OpenStudy (anonymous):

what is logbasex27=-3 explain please

OpenStudy (anonymous):

\[\log_{x} 27=-3\]

OpenStudy (anonymous):

\[x^{-3}=27\] is a start

OpenStudy (ash2326):

We have \[ \log_{x} 27=-3\] we know that \[ \log_{a} b= c=> b=a^c\] so here we have \[ \log_{x} 27=-3\] \[=> 27= x^{-3}\] or \[ 27=(\frac{1}{x})^3\] We know that3^3=27 so \[\frac{1}{x}=3\] we get \[x=\frac{1}{3}\]

OpenStudy (bahrom7893):

i read the question wrong..

OpenStudy (anonymous):

\[x^{-3}=3^3\] \[x^{-3}=(\frac{1}{3})^{-3}\]

OpenStudy (anonymous):

oh i see now thanks for showing the steps